The length of a rectangle is twice it's width. The perimeter is 54cm
Find the length
Find the width

Answer by macston(5194)

W=width; L=length=2W Perimeter=P=2(L+W)
54 cm=2(L+W) Substitute for L
54 cm=2(2W+W)
54 cm=2(3W)
54 cm=6W Divide each side by 6.
9 cm=W ANSWER 1: Width is 9 cm.
L=2W=2(9 cm)=18 cm ANSWER 2: Length is 18 cm.
CHECK:
P=2(L+W)
54 cm=2(18 cm + 9 cm)
54 cm= 2(27 cm)
54 cm=54 cm